# Down with the Hierarchy: The ‘H’ in HNSW Stands for “Hubs”

### v1

@@ 4. Reproduction of Prior Studies
In this section, we present two replicability studies of prior experimental evaluations on the efficacy of HNSW versus a non-hierarchical NSW index. As we discussed in the previous section, these prior works possess limitations in experimental design, scope of benchmarking datasets, and a lack of analysis into understanding the results, which motivates our work in this paper. Nevertheless, we use these prior studies as a starting point to see if we can independently replicate these results via our own FlatNav implementation. Such a reproduction would both further validate the soundness of these previous experiments over the test of time as well as provide validation of the correctness of FlatNav before we proceed to new, larger-scale benchmarks.
4.1. Reproduction of Malkov and Yashunin (2016)
For our first replicability study, we look to reproduce one of the core experimental result of the original HNSW paper (Malkov and Yashunin, 2016). To validate that the hierarchical layers provide a speedup over a flat NSW index, Malkov and Yashunin benchmark their HNSW implementation (available through the hnswlib library) against the sw-graph NSW implementation. Furthermore, we also add our own flat graph implementation via our flatnav library as an additional comparison. Following the original HNSW paper, we generate two synthetic random hypercube datasets consisting of 4 and 8 dimensions, respectively. In particular, each dimension of these random vectors is sampled uniformly at random from $[0,1)$. We report the results of this replicability study in Figure 2 where we find that HNSW indeed achieves superior latency than flatnav and sw-graph on these low-dimensional datasets.
@@ 4.1. Reproduction of Malkov and Yashunin (2016)
Figure 2. Mean Latency vs. Recall of HNSW, SW-Graph and FlatNav across dimensions $d=4,8$. We find that HNSW achieves lower mean latency than its flat index counterparts in FlatNav and SW-Graph, matching the findings of Malkov and Yashunin (2016)
4.2. Reproduction of Lin and Zhao (2019)
We now present our results reproducing the findings of Lin and Zhao (2019) in their 2019 benchmarking study, which concluded that the hierarchical structure of HSNW is only beneficial for low-dimensional data. Although Lin and Zhao primarily focus on synthetic, randomly generated datasets, we looked to see if we could independently replicate these findings before turning our attention to benchmarking on modern, large-scale embedding datasets in the next section.
Following the same setup as (Lin and Zhao, 2019), we generate a series of random vector datasets of varying dimensionality where each vector component is sampled uniformly at random from the range $[0,1)$. In particular, we consider dimensionalities of $d=4,8,16$ and 32. As in (Lin and Zhao, 2019), we set the number of near neighbors to retrieve to $k=1$ (departing from the default of $k=100$ we use elsewhere in this paper).
Figure 3. Median Latency vs. Recall of HNSW and FlatNav across dimensions $d=4,8,16,32$. We observe that the hierarchical structure accelerates search only when $d<32$, matching the findings of Lin and Zhao (2019)
In summary, we successfully replicated the experiments benchmarking HNSW versus a flat graph from two prior research papers. Notably, both of these previous works primarily experiment with randomly generated vector data with very low dimensionality by the standards of modern machine learning. Coupled with our findings in the next section where we find no discernible difference between HSNW and a flat graph index on high-dimensional datasets, our results suggest a simple decision criterion for selecting a search index: For dimensionality $d<32$, HNSW and the hierarchy provide a speedup. Otherwise, the simplicity and memory savings of a flat index provide more benefit.
@@ 5. FlatNav Benchmarking Experiments
A challenge in designing such benchmarking experiments is separating the confounding impact of performance engineering in the implementation of an approach from any purely algorithmic advantages of a method. We argue that this is a pitfall that has clouded several past benchmarking studies in ANN search. To circumvent this challenge, we fix the implementation in our experimental design such that the same code is used in constructing the base navigable small world graph. In particular, we use the popular hnswlib library as our baseline HNSW implementation. To benchmark the flat index performance, we extract the bottom graph layer from hnswlib and reimplement the HNSW’s greedy search heuristic over this flat graph. To the best of our knowledge, we provide the first benchmarking study on the utility of HNSW’s hierarchy that makes an effort to achieve parity in the implementations of the two techniques to center the focus on the central algorithmic question.
@@ 5.2. Datasets and Compute
Dataset Dimensionality Collection Size Num Queries
@@
Table 1. Dataset Statistics. The datasets marked by $\dagger$ are from the BigANN benchmarks (Simhadri et al., 2022). The remaining are taken from ANN Benchmarks (Aumüller et al., 2018)
@@ 5.3. Latency Results
Figure 4. p50 Latency vs. Recall for HNSW and FlatNav over four 100M-scale benchmark datasets with dimensionality between 96 and 200. We observe that the flat index achieves essentially identical performance to the hierarchical HNSW index.
@@
Figure 6. p50 Latency vs. Recall for HNSW and FlatNav over five ANN Benchmark datasets ranging in dimensionality from 96 to 960. The flat index again achieves practically indistinguishable performance from the hierarchical graph index.
@@
Figure 8. p50 Latency vs. Recall for HNSW and FlatNav over the GloVe ANN Benchmark datasets with respective dimensionalities 25, 50, 100, and 200. Again, we see no clear difference between the two indexes in terms of performance.
@@ 5.4. Memory
In this section we measure the memory savings from removing the hierarchy by running a memory profiler during index construction. In terms of memory allocation, similarly to flatnav, hnswlib allocates static memory during index construction comprising base layer node allocation, a visited node list and a list of mutexes in a multi-threaded setting. Additionally, it also incurs memory cost attributable to the hierarchy, particularly maintaning the dynamically allocated links between nodes at each layer.
We benchmarked both libraries against the BigANN 100M dataset shown in Table 1 and traced memory consumption during the index construction stage parallelized over $32$ cores. We observe that, at its peak, hnswlib reaches $183$ GBs of RAM while flatnav peaks at $113$ GBs, demonstrating up to $38\%$ reduction in peak memory consumption. This shows that we are able to save significant memory by removing the hierarchy, and it is likely that we can optimize flatnav implementation to save memory further.
@@ 6. The Hub Highway Hypothesis
We remark that the existence of hub nodes in high-dimensional space is not a new observation and has in fact been studied in the prior literature on small world graphs. The novelty of our hypothesis lies in connecting the idea of hubness to the notion of accelerating near neighbor search in proximity graphs. In particular, we conjecture that near neighbor queries over proximity graphs in high dimensions often spend the majority of their time visiting hub nodes early on in the search process before converging to a local neighborhood of near neighbors. This procedure succeeds because hub nodes are very well connected to other parts of the graph and thereby efficiently route queries to the appropriate neighborhood in much the same manner that the layered hierarchy purports to do.
In many applications, such as clustering, hubness is often considered an undesirable property that harms algorithmic performance, thereby motivating the need for hubness reduction techniques (Zelnik-Manor and Perona, 2004). On the contrary, our Hub Highway Hypothesis presents a case where hubness offers a distinct advantage in a practical algorithmic setting.
@@ 6.1. Methodology
In this section, we will present a formal definition and characterization of hubness and present empirical results in support of our hypothesis.
@@
Dataset Dimensionality Collection Size Num Queries
@@
Synthetic and ANN Benchmark Datasets: To illustrate evidence for the hub-highway hypothesis, we experiment with both real and synthetic datasets as shown in Table 2. In addition to a subset of ANN datasets, we generate synthetic datasets by drawing vectors from the standard normal distribution. The choice of the normal distribution is important because, as (Low et al., 2013) observe, data sampled from the normal distribution exhibit a stronger hubness phenomenon due to strong density gradient than other distributions, such as sampling uniformly from a $d$-dimensional hypercube. Furthermore, since the hubness phenomenon is accentuated in high-dimensional spaces, we generate random data with increasing dimensionality which allows us to study the effect of these intrinsic highway structures that form as $d$ increases.
In the following three experiments we show empirical evidence supporting the hub-highway hypothesis. We start by examining the distribution of the number of times each node in the similarity search index is visited during search given a fixed number of queries, which we denote as $P_{m}(\mathbb{x}_{i})$ to indicate that node $\mathbb{x}_{i}$ is visited exactly $m$ times during a fixed number of $k$-NN queries for each one of the datasets in Table 2. We show that this distribution is skewed to the right, thus confirming that certain nodes (highway nodes) are visited a disproportionate number of times. Next, using this distribution, we selectively choose the most visited nodes in the similarity search graph to be the hub node clusters and show evidence that these nodes are more connected to each other than random nodes using hypothesis tests. In the last experiment, we show that not only are these highway nodes more connected, they also allow for faster graph traversals for queries, hence supporting the claim that we no longer need hierarchy in high dimensional vector search where fast query times evolve as a result of the presence of the hub-highways.
@@ 6.2. Skewness of the Node Access Distribution
As described above, for this study we consider the discrete distribution $P_{m}(\mathbb{x}_{i})$, $\mathbb{x}_{i}\in\mathcal{D}$, where $\mathcal{D}$ is one of the datasets in Table 2. For all these experiments we construct a HNSW index with parameters listed in Table 3 and import the base layer of the graph into FlatNav to construct a flat index.
@@
(a) Cosine KDE Plot
(b) L2 KDE Plot
Figure 9. Log-normalized node access count distribution $P_{m}(\mathbb{x}_{i})$ plotted with kernel smoothing (Kernel Density Estimation). (a) Distribution for different datasets that use cosine similarity along with the skewness $\tilde{\mu}_{3}$ of $P_{m}(\mathbb{x}_{i})$. (b) Distribution for the datasets that use $l_{2}$ distance and the corresponding skweness values.
Figure 9 shows the log-normalized node access count distribution for different datasets. We observe that as the dimension $d$ increases, for $\ell_{2}$ distance-based datasets this distribution becomes more skewed to the right, indicating that a subset of nodes are visited more often than the rest of the nodes in the graph. It is worth noticing that even for $d\in\{1024,1536\}$, the cosine distance function exhibits anti-hub properties, effectively forcing little to no hubness in the data. This is consistent with the results observed by (Radovanovic et al., 2010). This increased skewness of the $P_{m}(\mathbb{x}_{i})$ distribution as $d$ increases explains how highway nodes occur as a result of the intrinsic dimensionality of the datasets.
@@ 6.3. Subgraph Connectivity of the Hub-Highway Nodes
Having shown how highway nodes occur, in the next set of experiments we show empirical evidence confirming that these nodes exhibit stronger connectivity in the graph than any given random nodes in the similarity search graph. We set up a hypothesis test as follows with the null hypothesis that hub nodes are not any more connected to each other than randomly chosen nodes. Let $\mathbb{x}_{i},1\leq i\leq n$, be the vectors in a similarity search index for any dataset $\mathcal{D}$ under consideration.
@@
  • For each identified hub node in the index, we compute the number of hub nodes adjacent to it. This yields a discrete distribution of the number of hub nodes to which each hub node is connected.
@@
  • Similarly, we select a set of random nodes in the set $V\coloneqq\mathcal{D}\setminus\bigcup_{i=1}^{n}h(\mathbb{x}_{i})=1$. For each node $\mathbb{x}_{i}\in V$, we compute the number of hub nodes adjacent to it to get an equivalent distribution.
@@
We set up both a parametric test with a two-sample t-test and a non-parametric test using the Mann-Whitney U-test (Mann and Whitney, 1947), which is suitable particularly for the ANN datasets since it makes no normality assumption. A well-known consequence of the Central Limit Theorem (CLT) (Durrett, 2019) is that in non-parametric statistical tests like the Mann-Whitney U test, as the sample size increases, the U-statistic converges to a normal distribution. This, in turn, leads to $p$-values that approach 0 whenever there is a true difference between any two populations under consideration. To avoid this pitfall of the $p$-value, we also consider effect size in our test and constrain the sample size to be 1000 for both populations.
Table 4 shows the $p$-value results in the regime where we use the p95 of the node access distribution to select hub nodes. Assuming the regular 0.05 threshold for rejecting the null hypothesis, notice that in all cases but 5 we reject the null hypothesis under the Mann-Whitney $U$-test (and all cases but 6 for the two-sample $t$-test). Unsurprisingly, for the synthetic datasets, the effect size is much higher for $l$-2 datasets since this distance metric induces stronger hubness in the dataset, effectively increasing the likelihood that the subgraph induced by hub nodes is more connected. It is worth noting that the Yandex-DEEP benchmark does not exhibit a similar hub node subgraph connectivity effect seen in the majority of the other benchmarks. Yandex-DEEP embeddings were generated from the last fully-connected layer of GoogleNet (Szegedy et al., 2014) by applying PCA and $l$-2 normalization. In this case $l$-2 normalization alleviates the effects of hubs since the magnitude of the vectors does not disproportionately affect distance computations.
Similarly, Table 5 shows the $p$-value results in the case we use the p99 threshold of the node access counts to select hub nodes. With the 0.05 statistical significance threshold we are able to rejec the null hypothesis in all cases but 1 for both the Mann-Whitney $U$-test and the two-sample $t$-test. In this case we realize much higher effect sizes, which is not surprising since the p99 threshold selects an even more refined set of hub nodes. Thus, the two statistical tests show strong evidence that hub nodes form more connected subgraphs than those induced by randomly sampling nodes in the similarity search graph.
@@
Table 4. Two-sample T-test and Mann-Whitney U-test p-values. Hub nodes are selected using the p95 threshold for the node access distribution.
@@
Table 5. Two-sample T-test and Mann-Whitney U-test p-values. Hub nodes are selected using the p99 threshold for the node access distribution.
@@
6.4. Hub-Highway Nodes Allow Queries to Traverse the Graph Faster
In this section we show that not only are highway nodes more connected to each other than random nodes, they also enable fast traversals along the similarity search graph. Intuitively this is not surprising given our results that these nodes are more connected to each other and have long range connections between them that allow queries to navigate across distant regions in navigable small world graphs efficiently. To measure this effect, we construct similarity search indexes and for each query $q$, track the sequence of nodes visited before beam search terminates.
Suppose $\mathbb{x}_{1},\mathbb{x}_{2},\ldots,\mathbb{x}_{l}$ is a length-$l$ sequence of such nodes for query $q.$ We use the hub node assignment heuristic discussed in section 6.3 to get a label $h(\mathbb{x}_{i})$ for each node in the sequence. We then split the sequence into bins and compute the percentage of hub nodes visited in each bin. For bin $B_{i}$, this is given by

### v2

@@ 4. Reproduction of Prior Studies
In this section, we present a replicability study using four flatnav NSW implementation. In particular, we revisit the experimental design of two prior works in the literature: the original 2016 HNSW paper of Malkov and Yashunin (2016) and a subsequent 2019 paper from Lin and Zhao (2019) that found limitations with the hierarchical component of HNSW. As we discussed in the previous section, these prior works possess limitations in experimental design, scope of benchmarking datasets, and a lack of analysis into understanding the results, which motivates our work in this paper. Nevertheless, we use these prior studies as a starting point to see if we can independently replicate these results via our own FlatNav implementation. Such a reproduction would both further validate the soundness of these previous experiments over the test of time as well as provide confirmation of the correctness of FlatNav before we proceed to new, larger-scale benchmarks.
Following the same setups as Malkov and Yashunin (2016) and Lin and Zhao (2019) we generate a series of random vector datasets of varying dimensionality where each vector component is sampled uniformly at random from the range $[0,1)$. In particular, we consider dimensionalities of $d=4,8,16$ and 32. As in (Lin and Zhao, 2019), we set the number of near neighbors to retrieve to $k=1$ (departing from the default of $k=100$ we use elsewhere in this paper). We also tried including the sw-graph NSW baseline (Boytsov and Naidan, 2013) that Malkov and Yashunin (2016) used in their evaluation to benchmark against HNSW, but we were not able to run this older library successfully. However, we were able to replicate these prior findings using our own flatnav implementation which is conceptually identical to sw-graph but with more software optimizations to achieve engineering parity with hnswlib (detailed in Section 5.1).
Figure 3. Median Latency vs. Recall of HNSW and FlatNav across dimensions $d=4,8,16,32$. We observe that the hierarchical structure accelerates search only when $d<32$, matching the findings of Lin and Zhao (2019). Our results demonstrating a significant advantage with HNSW on synthetic datasets with dimensionality $d=4$ and $d=8$ also match the findings of the original HNSW paper Malkov and Yashunin (2016)
As shown in Figure 3, we successfully replicated the experiments benchmarking HNSW versus a flat NSW graph from two prior research papers. Notably, both of these previous works primarily experiment with randomly generated vector data with very low dimensionality by the standards of modern machine learning. Coupled with our findings in the next section where we find no discernible difference between HSNW and a flat graph index on high-dimensional datasets, our results suggest a simple decision criterion for selecting a search index: For dimensionality $d<32$, HNSW and the hierarchy provide a speedup. Otherwise, the simplicity and memory savings of a flat NSW index provide more benefit.
@@ 5. FlatNav Benchmarking Experiments
A challenge in designing such benchmarking experiments is separating the confounding impact of performance engineering in the implementation of an approach from any purely algorithmic advantages of a method. We argue that this is a pitfall that has clouded several past benchmarking studies in ANN search. To circumvent this challenge, we fix the implementation in our experimental design such that the same code is used in constructing the base navigable small world graph. In particular, we use the popular hnswlib library as our baseline HNSW implementation. To benchmark the flat NSW index performance, we extract the bottom graph layer from hnswlib and reimplement the HNSW’s greedy search heuristic over this flat graph. To the best of our knowledge, we provide the first benchmarking study on the utility of HNSW’s hierarchy that makes an effort to achieve parity in the implementations of the two techniques to center the focus on the central algorithmic question.
@@ 5.2. Datasets and Compute
Dataset Dimensionality # Points # Queries
@@
Table 1. Dataset Statistics. The datasets marked by $\dagger$ are from the BigANN benchmarks (Simhadri et al., 2022). The remaining are taken from ANN Benchmarks (Aumüller et al., 2018).
@@ 5.3. Latency Results
Figure 4. p50 Latency vs. Recall for HNSW and FlatNav over four 100M-scale benchmark datasets with dimensionality between 96 and 200. We observe that the flat NSW index achieves essentially identical performance to the hierarchical HNSW index.
@@
Figure 6. p50 Latency vs. Recall for HNSW and FlatNav over five ANN Benchmark datasets ranging in dimensionality from 96 to 960. The flat NSW index again achieves practically indistinguishable performance from the hierarchical graph index.
@@
Figure 8. p50 and p99 Latency vs. Recall for HNSW and FlatNav over the GloVe ANN Benchmark datasets with respective dimensionalities 25, 50, 100, and 200. We observe that HNSW has a minor advantage in low dimensions and that FlatNav has a minor advantage in high dimensions, but also that there is no clear difference in performance.
@@ 5.4. Memory
Dataset # Data hnswlib Memory flatnav Memory
BigANN 100M 183 113
Microsoft SpaceV 100M 104 85.5
Yandex DEEP 100M 100 60.7
Table 2. Peak Index Construction Memory in GBs. We observe that flatnav requires considerably less memory during construction compared to hnswlib.
In this section we measure the memory savings from removing the hierarchy by running a memory profiler during index construction. In terms of memory allocation, similarly to flatnav, hnswlib allocates static memory during index construction comprising base layer node allocation, a visited node list and a list of mutexes in the multi-threaded setting. Additionally, it also incurs memory cost attributable to the hierarchy, particularly maintaning the dynamically allocated links between nodes at each layer.
We benchmarked both libraries against a subset of the BigANN benchmarks. Table 2 shows the peak memory allocated by the two implementations during index construction for the BigANN, Microsoft SpaceV and Yandex DEEP benchmarks. Since multithreading has a runtime overhead, we fix the number of cores to 32 in each one of the stated benchmark. For BigANN, we observe a $38\%$ reduction in peak memory, a $39\%$ reduction for Yandex DEEP, and an $18\%$ reduction for the Microsoft SpaceV benchmark. This shows that we are able to save significant memory by removing the hierarchy, and it is likely that we can optimize flatnav implementation to save memory further.
@@ 6. The Hub Highway Hypothesis
We remark that the existence of hub nodes in high-dimensional space is not a new observation (Radovanovic et al., 2010). The novelty of our hypothesis lies in connecting the idea of hubness to the notion of accelerating near neighbor search in ANN proximity graphs. In particular, we conjecture that near neighbor queries over proximity graphs in high dimensions often spend the majority of their time visiting hub nodes early on in the search process before converging to a local neighborhood of near neighbors. This procedure succeeds because hub nodes are very well connected to other parts of the graph and thereby efficiently route queries to the appropriate neighborhood in much the same manner that the layered hierarchy purports to do.
In many applications, such as clustering, hubness is considered an undesirable property that harms algorithmic performance, motivating the need for hubness reduction techniques (Zelnik-Manor and Perona, 2004). On the contrary, our Hub Highway Hypothesis presents a case where hubness offers a distinct advantage in a practical algorithmic setting.
@@
6.1. Intuition:
The formation of a hub highway can be understood through concentration of measure phenomena.
Concentration of distances: Under many data-generating distributions, the expected distances between randomly-drawn points converge to a constant value as dimension increases (Beyer et al., 1999). This suggests that in sufficiently high-dimensional spaces, we do not need to explicitly encourage the formation of “long-range” connections. Since all distances are approximately equal, all connections have similar length and a few edge traversals should span the graph of a high-dimensional dataset.
Concentration of measure: A well-known fact about high-dimensional geometry is that volume concentrates at the extrema of a shape (e.g., the surface of a ball or the faces of a hypercube). These portions of the distribution contain a disproportionate number of points and frequently contain the hubs of the metric space. For example, Low et. al. demonstrate that the the hubs of a uniform hypercube distribution are located at the corners (Low et al., 2013), which are exactly the points we would wish to connect via long-range connections in a hierarchical graph.
@@ 6.2. Methodology
Argument sketch: We will demonstrate the Hub Highway Hypothesis by providing empirical evidence for the following claims.
  1. (1)
    Some nodes are visited by queries much more frequently than others. The relative popularity of these hub nodes is explained by the hubness phenomena that arises in high-dimensional datasets.
  2. (2)
    The hub nodes tend to be connected to each other, forming a well-connected subgraph of hubs.
  3. (3)
    Queries visit many hub nodes early in the search process, before visiting less well-traversed neighborhoods.
Empirical measures of hubness: To support the first part of our argument, we will need a formal definition and characterization of hubness.
@@
Dataset Dimensionality # Data # Queries
@@
Synthetic and ANN Benchmark Datasets: To illustrate evidence for the Hub Highway Hypothesis, we experiment with both real and synthetic datasets as shown in Table 3. In addition to a subset of ANN Benchmark datasets, we generate synthetic datasets by drawing vectors from the standard normal distribution. The choice of the normal distribution is important because, as (Low et al., 2013) observe, data sampled from the normal distribution exhibit a stronger hubness phenomenon due to strong density gradient than other distributions, such as sampling uniformly from a $d$-dimensional hypercube. Furthermore, since the hubness phenomenon is accentuated in high-dimensional spaces, we generate random data with increasing dimensionality which allows us to study the effect of these intrinsic highway structures that form as $d$ increases.
In the following three experiments we show empirical evidence supporting the Hub Highway Hypothesis. We start by examining the distribution of the number of times each node in the similarity search index is visited during search given a fixed number of queries, which we denote by $P_{m}(\mathbb{x}_{i})$ to indicate that node $\mathbb{x}_{i}$ is visited exactly $m$ times during a fixed number of $k$-NN queries for each one of the datasets in Table 3. We show that this distribution is skewed to the right, thus confirming that certain nodes (highway nodes) are visited a disproportionate number of times. Next, using this distribution, we selectively choose the most visited nodes in the similarity search graph to be the hub node clusters and show evidence that these nodes are more connected to each other than random nodes using hypothesis tests. In the last experiment, we show that not only are these highway nodes more connected, they also allow for faster graph traversals for queries, hence supporting the claim that we no longer need hierarchy in high dimensional vector search where fast query times evolve as a result of the presence of the hub-highways.
@@ 6.3. Skewness of the Node Access Distribution
As described above, for this study we consider the discrete distribution $P_{m}(\mathbb{x}_{i})$, $\mathbb{x}_{i}\in\mathcal{D}$, where $\mathcal{D}$ is one of the datasets in Table 3. For all these experiments we construct a HNSW index with parameters listed in Table 4 and import the base layer of the graph into FlatNav to construct a flat NSW index.
@@
Figure 9. Log-normalized node access count distribution $P_{m}(\mathbb{x}_{i})$ for datasets using angular distance (left) and $\ell_{2}$ distance (right). Note that the x-axis is also log-normalized, so all histograms are skewed to the left – but the $\ell_{2}$ datasets have a much greater skew.
Figure 9 shows the log-normalized node access count distribution for different datasets. We observe that as the dimension $d$ increases, this distribution becomes more skewed to the right for $\ell_{2}$ distance-based datasets, indicating that a subset of nodes are visited much more often than the rest of the nodes in the graph. The cosine distance, on the other hand, is known to have anti-hub properties that prevent such a dramatic skew even for $d\in\{1024,1536\}$. This is consistent with the results observed by (Radovanovic et al., 2010). The increased skewness of the $P_{m}(\mathbb{x}_{i})$ distribution as $d$ increases suggests that highway nodes occur as a result of the intrinsic dimensionality of the datasets.
Confounding mechanisms: While Figure 9 confirms the existence of hub nodes in the dataset, it does not distinguish between hubs that arise from the properties of the underlying metric space and hubs that form through some other mechanism. It is possible that preferential attachment explains the formation of hubs, since NSW graphs are built incrementally by sequentially adding points to an existing graph. Nodes that are added early in graph construction may become hubs by accumulating a greater-than-average share of inbound graph links, rather than by being popular neighbors in the metric space.
To investigate the effects of preferential attachment, we computed the variance ($R^{2}$ of a linear model) in the empirical node access distribution explained by the insertion ordering(Table 5). We log-transformed both the node access count and the insertion order before running the linear model and confirmed that the residuals are approximately normal by visually examining the QQ plots ($p<10^{-6}$ for all models).
We observe a modest effect from the node insertion order in our synthetic data. This is particularly true for the angular datasets, which we believe to be due to the weaker hubness phenomena produced by the angular distance metric. Preferential attachment may account for a relatively greater share of the node access distribution when metric hubs are not present to heavily skew the distribution.
Ideally, we would repeat this analysis using the $K$-occurrence distribution, to show that the hubness of the metric space is more strongly predictive of the node acccess count than the insertion order. Unfortunately, it is not feasible to compute the $K$-occurrence distribution due to the $O(n^{2})$ brute-force computation cost. However, we believe that the results in Table 5 still support the idea that the dimensionality of the metric space strongly contributes to the formation of hubs in the Hub-Highway Hypothesis, especially when combined with the evidence in Figure 9.
Table 5. Correlation analysis for node insertion order and node access count, to determine effects of preferential attachment.
Dataset Dimension Explained Variance (%) IID Normal (Angular) 16 $3.0$% IID Normal (Angular) 32 $8.7$% IID Normal (Angular) 64 $16.6$% IID Normal (Angular) 128 $23.3$% IID Normal (Angular) 256 $24.6$% IID Normal (Angular) 1024 $24.1$% IID Normal (Angular) 1536 $23.9$% IID Normal (L2) 16 $<0.1$% IID Normal (L2) 32 $3.1$% IID Normal (L2) 64 $7.6$% IID Normal (L2) 128 $8.7$% IID Normal (L2) 256 $7.1$% IID Normal (L2) 1024 $7.1$% IID Normal (L2) 1536 $6.6$% GloVe (Angular) 100 $0.2$% NYTimes (Angular) 256 $0.3$% GIST (L2) 960 $<0.1$% Yandex-DEEP (L2) 96 $0.3$% Microsoft-SpaceV (L2) 100 $0.5$%
@@ 6.4. Subgraph Connectivity of the Hub-Highway Nodes
Having demonstrated the occurrence of highway nodes, in the next set of experiments we present empirical evidence confirming that these nodes exhibit strong connectivity in the graph. One reasonable approach could be to identify hub nodes by thresholding the node access distribution (e.g., to the top 5% by access frequency) and determining whether these nodes form a connected component. However, this is strongly influenced by the threshold probability, and in any case the hubs need not form a single subgraph to support the Hub Highway Hypothesis.
Instead, we rely on statistical hypothesis testing to determine whether hub nodes are likely to connect to other hub nodes. We begin by explaining our procedure to identify hubs. Let $\mathbb{x}_{i},1\leq i\leq n$, be the vectors in a similarity search index for any dataset $\mathcal{D}$ under consideration.
@@
  • We wish to estimate the likelihood that a randomly-chosen out-neighbor of a hub node is, itself, a hub. To do so, we examine the 1-hop out-degree expansion of each hub and count the number of adjacent hub nodes. By repeating this process for each hub node, we obtain a discrete distribution for the number of hubs to which each hub node is connected.
@@
  • Similarly, we select a set of random non-hub nodes from $V\coloneqq\mathcal{D}\setminus\bigcup_{i=1}^{n}h(\mathbb{x}_{i})=1$. For each node $\mathbb{x}_{i}\in V$, we compute the same quantity to find the number of hubs with which non-hubs are connected, allowing us to construct the equivalent distribution for non-hubs.
Our hypothesis test investigates whether these two distributions differ. The null hypothesis is that hubs and non-hubs are equally likely to attach to hubs, and the alternative hypothesis is that hubs display different connectivity behavior than the rest of the graph. A finding that hubs are statistically more likely to attach to other hubs will support the Hub Highway Hypothesis.
@@
We measure statistical significance using the standard two-sample parametric $t$-test as well as the non-parametric Mann-Whitney U-test (Mann and Whitney, 1947), which is particularly suitable for ANN datasets because it does not require the normality assumption. Note that we consider dataset sizes larger than those often encountered in hypothesis testing ($n>10^{3}$). A well-known consequence of the Central Limit Theorem (CLT) (Durrett, 2019) is that the U-statistic used in non-parametric tests such as the Mann-Whitney test will converge to a normal distribution as the sample size increases. This, in turn, leads to $p$-values that approach 0 whenever there is a true difference between any two populations under consideration. For this reason, we also consider the effect size in our test and constrain the sample size to be 1000 for both populations.
Table 6 shows the results when we use the top-5% threshold to select hub nodes. Assuming the regular 0.05 threshold for rejecting the null hypothesis, note that in all by five cases we reject the null hypothesis under the Mann-Whitney $U$-test (and all by six cases using the two-sample $t$-test). Unsurprisingly, the effect size is much higher in the synthetic $\ell_{2}$ datasets, likely because this metric and distribution induce strong metric hubness properties in the dataset and thus a more connected subgraph.
Notably, the Yandex-DEEP benchmark does not exhibit the subgraph connectivity observed in most other benchmarks. Unlike the other embeddings in our collection, Yandex-DEEP was obtained by applying PCA and $\ell_{2}$ normalization to the last fully-connected layer of GoogleNet (Szegedy et al., 2014). Theoretically, $\ell_{2}$ normalization should reduce the prevalence of hubs by restricting the distribution to the surface of the unit sphere. Yandex-DEEP also has the most uniform node access distribution of the $\ell_{2}$ datasets in Figure 9. We believe that this may explain why Yandex-DEEP hubs are not as well-connected.
Table 7 shows the results when hubs are identified using the top-1% threshold. Here, we are able to reject the null hypothesis ($p<0.05$) in all but one case for both the Mann-Whitney $U$-test and the two-sample $t$-test. We also observe much larger effect sizes with the top-1% selection criterion, demonstrating that the strongest hubs cluster tightly in the graph. Taken together, these two statistical tests support the idea that hub nodes are often connected to other hubs.
@@
Table 6. Two-sample $t$-test and Mann-Whitney U-test results. Hub nodes are selected using the P95 threshold of the node access distribution.
@@
Table 7. Two-sample $t$-test and Mann-Whitney U-test results. Hub nodes are selected using the P99 threshold of the node access distribution.
@@
6.5. Hub-Highway Nodes Enable Fast Traversal
In this section, we show that not only are highway nodes strongly connected to each other, they also allow queries to quickly traverse the similarity search graph. While it is not surprising that a well-connected subgraph of frequently visited nodes would enable this behavior, it is not immediately clear that queries use the highway in the way predicted by our hypothesis – to quickly identify a neighborhood for deep exploration. To investigate this question, we track the sequence of nodes visited during beam search. We examine thousands of queries to determine the fraction of time spent on hub nodes in different phases of search.
Since beam search takes a variable number of steps for each query, we normalize by the total search length when presenting the results. More formally, suppose that $\mathbb{x}_{1},\mathbb{x}_{2},\ldots,\mathbb{x}_{l}$ is a length-$l$ sequence of such nodes visited by a query. We use the hub node assignment heuristic discussed in section 6.4 to label $h(\mathbb{x}_{i})$ each of these nodes as hubs / non-hubs. Then, we then split the sequence into bins and compute the prevalence of hubs in each bin. For bin $B_{i}$, this is given by